# Scalable architecture for trapped-ion quantum computing using RF traps and dynamic optical potentials

### 08 Jan 2024 v2

@@ VI Parallel multi-qubit logic and crosstalk mitigation
In trapped-ion based quantum computers, logical gates between qubits, i.e. entanglement, is typically facilitated by spin-dependent forces acting on normal-modes of motion of the ion-crystal [61, 62]. Thus adequate control of the normal-mode structure is crucial for scaling-up the trapped-ion quantum register size. Specifically, utilizing many modes of motion simultaneously enables programmable long-range interactions [12, 13, 14], which are significant for efficient implementation of various quantum computational tasks [63, 64, 65, 66, 53]. However utilizing multiple modes of motion becomes increasingly challenging in large ion-crystals due to the overwhelming complexity of the required qubit drive, which comes about as a NP-hard optimization problem [12], as well as an apparent slow-down of the feasible gate duration [14] which, in absence of segmentation, scales as $N^{2}$. Thus segmentation plays here a crucial role as well in generating independent large, yet manageable, computational qubit cells.
The heuristic estimation for crosstalk error, provided by $\varepsilon_{\text{BW,b}}$ above, can be improved by a detailed consideration of the transverse mode-structure. Specifically, based on the quadratic-form of qubit-qubit entanglement discussed thoroughly in [14], the coupling term between ions $c$ ($c^{\prime}$) in cells $s$ ($s^{\prime}$), mediated by band $b$, is estimated as
@@
A few percent crosstalk level, reflected by both $\varepsilon_{\text{BW},b}$ and $\varepsilon_{J,b}$, marks an efficient decoupling between cells. Nevertheless, the remaining coupling is too high for high-fidelity quantum operations, and restricts performing simultaneous logical gate on distinct cells. However this crosstalk level is low enough such that it can be mitigated perturbatively in a scalable manner. Here we specifically rely on the large-scale fast (LSF) method, presented in Ref. [14].
@@
In essence, LSF makes use of a multi-tone drive in order to generate programmable multi-qubit gates, which naively involves solving a quadratically constrained optimization problem of dimension $N^{2}$, and generates the unitary gate, $\exp\left(\sum_{n,m=1}^{N}\varphi_{n,m}\sigma_{x}^{\left(n\right)}\sigma_{x}^{\left(m\right)}\right)$, with $\sigma_{x}^{\left(n\right)}$ a Pauli-$x$ operator acting on the $n$th qubit and $\varphi_{n,m}$ a target multi-qubit coupling matrix. The hardness of the optimization problem is mitigated in LSF with a key realization; namely, that for each fixed system, any non-vanishing drive solution to a special ‘zero-phase’ version of the optimization problem ($\varphi^{\left(0\right)}_{n,m}=0$) can be efficiently converted to a high-fidelity solution of any general problem. Finding such a zero-phase solution (ZPS) is generally a far easier task than directly solving the general problem.
The operational approach of LSF is naturally adopted to our architecture, making the design of multi-qubit entanglement, simultaneously in all cells, fast and scalable. Specifically, we use LSF in order to generate ZPSs for an effective ‘typical’ cell of the bulk of the ion-crystal (see the SM [47]). Up to small deviations that are corrected below, this ZPS is adequate to all of the ion-crystal’s cells. Using LSF, different multi-qubit gates are designed for each cell by converting the same ZPS. This classical computation can be performed fast and in parallel. After performing these steps we obtain a specific drive spectrum, for all computational and auxiliary qubits in the ion-crystal, that generates the required entanglement per-cell at once, $\exp\left(\sum_{s=1}^{S}\sum_{n,m=1}^{C}\varphi_{n,m}^{\left(s\right)}\sigma_{x}^{\left(n,s\right)}\sigma_{x}^{\left(m,s\right)}\right)$, with $\sigma_{x}^{\left(n,s\right)}$ a Pauli-$x$ operator acting on the $n$th qubit in cell $s$ and $\varphi_{n,m}^{\left(s\right)}$ the desired multi-qubit target of cell $s$.
@@
As discussed above, the unwanted coupling between cells incurs a $\sim 2.5\%$ crosstalk error, which is dominated by nearest-neighbor cells. Here this brings about a typical coupling of $\sim 0.025\varphi_{n,m}^{\left(s\right)}$ between ions in distinct cells. Therefore we make use of an additional mitigation layer on top of LSF. Crucially, the nearest-neighbor structure of the crosstalk enables our mitigation technique to scale favorably, in terms of classical computation resources, and does not impose constraints on the ion-crystal size.
@@
Specifically, we perform an iterative optimization by linearizing the quadratic constraints around the current solution, with the initial solution given by LSF as discussed above. The resulting linear equations provide conditions that resolve crosstalk as well as inaccuracies of the LSF solution that originate from the assumption of a typical cell. We do so locally, i.e. the linear conditions are formed for only two adjacent cells at a time, generating $C\left(C-1\right)$ linear constraints that improve the LSF results and $3C^{2}$ linear constraints that correspond to crosstalk terms. The iterative optimization is stopped when a target infidelity is reached. Crucially since the linear equations are local, relating only four adjacent cells at a time, they involve only $\mathcal{O}\left(C^{2}\right)$ linear constraints, allowing parallel optimization of the next four adjacent cells [47].
@@
Figure 7: Performance of the crosstalk mitigation method as a function of optical trapping potential $\omega_{\text{o.t.p}}$, evaluated on the entanglement used to implement a simulation of a 3D rectangular Ising model. Top: Error per cell, using Eq. (7), as a proxy for the entanglement infidelity of the entire multi-qubit operation. For $\omega_{\text{o.t.p}}\gtrsim\omega_{\text{rad}}$ the crosstalk compensated solution (dark-blue) shows a low, $<10^{-4}$, infidelity. As expected, the initial LSF solution, based on the effective typical cell (orange) shows a $\sim 10^{-2}$ infidelity. The mitigation degrades in the weak optical confinement regime, due to long-range coupling between cells that is not considered in the crosstalk mitigation method. Indeed this degradation appears when considering long-range cells (dashed purple), and not seen when only considering nearest-neighbor and onsite coupling (dashed pink). Additional infidelity due to residual coupling to motional modes (dashed light-blue) exhibits a similar behavior. Bottom: The required total drive coupling, given in units of the lowest transverse mode frequency (bottom). The crosstalk compensated solution (dark-blue) only incurs a small, $\sim 10\%$, overhead on the initial LSF solution (orange).
We study the efficacy of our crosstalk mitigation method by considering a segmented ion-crystal system used for a 3D Ising model simulation, as that shown in Fig. 2 above, with $S=5$, $d=4$ and $B_{A}=B_{B}=3$. We design the control pulses that implement the required qubit-qubit couplings, for various optical trapping frequencies. The resulting operation infidelity is evaluated, to leading order in the qubit-qubit coupling deviations and in residual ion displacements, as,
@@
Figure 7(top) presents the infidelity, $I$, of the initial LSF solution, based on the typical cell (orange) and the crosstalk corrected solution (dark-blue). It exhibits an apparent cross-over between two regimes at $\omega_{\text{o.t.p}}=2\nu$, at which point the transverse mode band structure is manifested. At low optical confinement the performance of uncorrected, typical-cell solutions (orange) is low, such that our leading-order estimation breaks-down. At strong optical confinement this solution generates an infidelity of $\sim 10^{-2}$, as expected from our analysis above. However the crosstalk-corrected solution (dark-blue) successfully generates solutions that have an infidelity that is better (lower) than $10^{-4}$ for the entire multi-qubit operation, throughout the strong optical confinement regime. We account for the different contributions to the infidelity of the crosstalk-corrected solution, given by Eq. (7). We note that $I_{\text{short}}$ (dashed pink), has a negligible contribution, regardless of optical confinement, while $I_{\text{long}}$ (dashed purple) reaches a low values only at the strong confinement regime. This behavior is expected as our optimization actively corrects for on-site and nearest-neighbor terms, and ignores other terms. We note that $I_{\text{long}}$ reaches a minimum at $\omega_{\text{o.t.p}}=2\nu$, and is slightly increased at higher optical trapping frequencies. This behavior is accounted by an effective screening of barrier ions that occurs when barrier bands are close to resonance with the bulk bands (see the SM [47]). Nevertheless, our analysis implies that strong optical confinement successfully decouples next nearest-neighbor (and further) cells, removing the need to take the entire ion-crystal into account, and enabling scalability. Similarly, due to a well-formed band structure, the motional infidelity, $I_{\text{motion}}$ (dashed light-blue), becomes small at strong optical confinement, and in general dominates the overall infidelity.
@@
We remark that elements of our architecture and crosstalk compensation method can also be implemented by global driving of entire cells using similar techniques to those in Refs. [12, 14].
@@ VII Mid-circuit measurements
In our optically segmented ion-crystal scheme, optically confined ions are well-suited for mid-circuit measurements, as their motion is separated to independent bands (top part of the spectra in Figs. 4 and 5), thus photon scattering heating remains local and does not heat the other crystal modes. Furthermore, the optical fields generating the trapping also substantially light-shift the atomic transition lines of the ions, specifically the photon emission lines, such that photon scattering is largely suppressed. Thus no physical shuttling of the ions into dedicated measurement regions is required, nor the use of multiple ion-species.
@@
Figure 8: Top: schematic of mid-circuit measurement protocol, showing three types of involved ions, computational qubit ions (top, purple circle), barrier ion, $B_{\text{B}}$, to be measured (middle, purple circle) and optically confined barrier ions, $B_{\text{A}}$ (bottom, purple full circle). Operations are ordered from left to right. Data is encoded on the computational qubit (e.g. $U_{s,n}^{\left(A\right)}$), and one of the qubit states of the ion is ‘shelved’ to a non-fluorescing state using local control. Next an intermediate optical segmentation configuration is used, confining both ions A and B barrier ions, and separating their motion from the computational qubits. The state of the barrier ions is measured and then reset (‘Prepare’). Lastly the optical segmentation configuration is switched. Bottom: Relevant atomic levels of ${}^{40}\text{Ca}^{+}$ ions, showing $4S_{1/2}$ the ground state manifold, the metastable $4D_{5/2}$ manifold and the short-lived $4P_{1/2}$ manifold. The $S\leftrightarrow P$ transition varies between the no-optical confinement (left) and optically confined (right) cases, allowing to separate the two cases spectrally (see main text). The main optical field used in the protocol are shown, namely at $729$ nm (red), coupling the $S\leftrightarrow D$ levels, at $400$ nm (dark-blue), generating Raman transition between qubit states as well as additional local light-shifts, and at $397$ nm (light-blue) coupling the $S\leftrightarrow P$ levels.
@@
We describe the protocol used to measure the qubit state of some target ions, without incurring decoherence on computational qubits, or heating of bulk modes. After the measurement is completed, we cool and re-initialize the measured ions. Figure 8 (top) illustrates these general steps involved in performing mid-circuit measurements. For simplicity we consider the required operations on three types of ions in the ion-crystal, i.e. computational qubit ions (top, purple circles), barrier ions to be measured (middle, purple circle), and optically confined barrier ions (bottom, purple full circles). Crucially, our measurement method can be performed in parallel throughout the entire ion-crystal on all required ions according to these roles. In our architecture ions may be optically confined specifically for the purpose of mid-circuit measurements, without having an essential role in segmentation of the ion-crystal, as showcased in the application of a QEC surface code, above.
@@
An additional locally-applied field resonant with the light-shifted $S\leftrightarrow P$ transition at $397$ nm is used (purple arrow). Since the $P$ level is broad, the required stability of the optically confining fields in order to maintain resonance is reasonable. The combination of local $400$ nm and $397$ nm fields, together with standard repump fields (not shown) accommodates for all the necessary steps - i.e., state detection, various cooling techniques, and state preparation. We note that the technical requirements for the $397$ nm field are less demanding than those of the optical confinement and gate drive fields at $400$ nm; the necessary power per ion is far lower, independent spectral control is not required, and the focal spot can generally be larger (by illuminating adjacent ions within the same beam).
@@
The mid-circuit measurement protocol shown in Fig. 8 is carried out by first encoding some information on ion B via the entanglement operations detailed above ( $U_{s,n}^{\left(A\right)}$), e.g. parity information of a plaquette shown in Fig. 3. Then, one of the qubit states of the ion to be measured is ‘shelved’ to the D manifold via a combination of $400$ nm light-shift and $729$ nm coupling, making this operation local on the ion.
Next, the segmentation configuration is changed to an intermediate setting where all barrier ions are illuminated. The purpose of optically confining the ion is to prepare it for being measured, by separating its motion from the computational qubits’ motion and to light-shift its $S\leftrightarrow P$ transition. We note that, for this protocol, the strength of the optical trapping potential should be adjusted so that all of the barrier bands are fully separated from the bulk bands (slightly above the value marked by the horizontal line in Fig. 4). Moreover, it is possible that the ion fully populates in the $D$ level, in which case it is not optically confined; thus existing confinement of other ions (purple full circle) cannot be relaxed. State detection is performed using the light-shifted $397$ nm field, followed by qubit reset and preparation using a combination of the $397$ nm and $400$ nm fields.
@@
During the detection and preparation steps, the optically confined ions scatter $397$ nm photons, such that they are required to be cooled down in the ‘prepare’ step, which can be done in parallel. The scattered photons are substantially detuned, by approximately $0.8$ GHz, from the absorption lines of the computational qubit ions, having a linewidth of few tens of MHz, thus negligibly influencing their state. Nevertheless the state of neighboring computational qubits, located several micrometers apart, may be further protected by shelving both of their qubit states to the $D$ level (at the shelving step). Alternatively they may simply remain unutilized.
Lastly the segmentation configuration can now be changed, exchanging the roles of computational and barrier ions. The cooled ions are returned to the bulk of the cells and in turn cool down the bulk modes via sympathetic cooling.

### 04 Nov 2024 v3

@@ VI Parallel multi-qubit logic and crosstalk mitigation
In trapped-ion based quantum computers, entanglement gates are typically facilitated by spin-dependent forces acting on normal-modes of motion of the ion-crystal [63, 57]. Thus adequate control of the normal-mode structure is crucial for scaling-up the trapped-ion quantum register size. Specifically, utilizing many modes of motion simultaneously enables programmable long-range interactions [13, 14, 15], which are significant for efficient implementation of various quantum computational tasks [64, 65, 66, 67, 54, 43, 68, 14]. However utilizing multiple modes of motion becomes increasingly challenging in large ion-crystals due to the overwhelming complexity of the required qubit drive, which comes about as a challenging optimization problem [13], as well as an apparent slow-down of the feasible gate duration [15] which, in absence of segmentation, scales as $N^{2}$. Thus segmentation plays here a crucial role as well in generating independent large, yet manageable, computational qubit cells.
The heuristic estimation for crosstalk error, provided by $\varepsilon_{\text{BW,b}}$ above, can be improved by a detailed consideration of the transverse mode-structure. Specifically, based on the quadratic-form of qubit-qubit entanglement discussed thoroughly in [15], the coupling term between ions $c$ ($c^{\prime}$) in cells $s$ ($s^{\prime}$), mediated by band $b$, is scaled by the factor,
@@
A few percent crosstalk level, reflected by both $\varepsilon_{\text{BW},b}$ and $\varepsilon_{J,b}$, marks an efficient decoupling between cells. Nevertheless, the remaining coupling is too high for high-fidelity quantum operations, and restricts performing simultaneous logical gates on distinct cells. However this crosstalk level is low enough such that it can be mitigated perturbatively in a scalable manner. Here we do so by relying on the large-scale fast (LSF) method, presented in Ref. [15].
@@
In essence, LSF makes use of a multi-tone drive in order to generate programmable multi-qubit gates, which naively involves solving a quadratically constrained optimization problem of dimension $N^{2}$, and generates unitary gates of the form $U=\exp\left(i\sum_{n,m=1}^{N}\varphi_{n,m}\sigma_{x}^{\left(n\right)}\sigma_{x}^{\left(m\right)}\right)$, with $\sigma_{x}^{\left(n\right)}$ a Pauli-$x$ operator acting on the $n$th qubit and $\varphi_{n,m}$ a target multi-qubit coupling matrix. This gate, together with arbitrary single qubit rotations realizes a universal gate set.
The operational approach of LSF is naturally adopted to our architecture, making the design of multi-qubit entanglement, simultaneously in all cells, fast and scalable. In particular, we use LSF for a ‘typical’ cell of the bulk of the ion crystal (see SM for details [48]) to obtain drive spectra corresponding to different target unitaries. These solutions could then, in principle, be applied to each cell within the ion crystal. Up to minor differences between the ‘typical’ and actual cells (to be corrected below), this would yield high-fidelity multi-qubit gates if applied on each cell separately.
@@
However, these operations are meant to run in parallel and, as discussed above, the unwanted coupling between cells incurs a $\sim 2.5\%$ crosstalk error - which is dominated by nearest-neighbor cells. Therefore we make use of an additional mitigation layer on top of LSF. Crucially, the nearest-neighbor structure of the crosstalk enables our mitigation technique to scale favorably, in terms of classical computation resources, and does not impose constraints on the ion-crystal size.
@@
Specifically, we perform an iterative optimization by linearizing the quadratic constraints associated with the target unitary around the current solution, with the initial solution given by LSF as discussed above. The resulting linear equations provide conditions that resolve crosstalk as well as inaccuracies of the LSF solution that originate from the assumption of a typical cell. We do so locally, i.e. the linear conditions are formed for only two adjacent cells at a time, generating $C\left(C-1\right)$ linear constraints that improve the LSF results and mitigate deviations of the cells from the ‘typical’ cell, and $3C^{2}$ linear constraints that mitigate crosstalk between cells. The iterative optimization is stopped when a target infidelity is reached. Crucially since the linear equations are local, relating only four adjacent cells at a time, they involve only $\mathcal{O}\left(C^{2}\right)$ linear constraints, allowing parallel optimization of the next four adjacent cells [48].
@@
Figure 7: Performance of the crosstalk mitigation method as a function of optical trapping potential $\omega_{\text{o.t.p}}$, evaluated on the entanglement used to implement a simulation of the 3D rectangular Ising model in Fig. 2. Top: Error per cell, using Eq. (7), as a proxy for the infidelity of the entire multi-qubit operation. For $\omega_{\text{o.t.p}}\gtrsim 2\nu$ the crosstalk compensated solution (dark-blue) shows a low, $<10^{-4}$, infidelity. As expected, the initial LSF solution, based on the effective typical cell (orange) shows a $\sim 10^{-2}$ infidelity. The mitigation degrades in the weak optical confinement regime, due to long-range coupling between cells that is not considered in the crosstalk mitigation method. Indeed this degradation appears when considering long-range cells (dashed purple), and not seen when only considering nearest-neighbor and onsite coupling (dashed pink). Additional infidelity due to residual coupling to motional modes (dashed light-blue) exhibits a similar behavior. Bottom: The required total drive coupling, given in units of the lowest transverse mode frequency (bottom). The crosstalk compensated solution (dark-blue) only incurs a small, $\sim 10\%$, overhead on the initial LSF solution (orange).
We study the efficacy of our crosstalk mitigation method by considering a segmented ion-crystal system used for a 3D Ising model simulation, as that shown in Fig. 2 above, with $S=5$, $d=4$ and $B_{A}=B_{B}=3$. We design the control pulses that implement the required qubit-qubit couplings, for various optical trapping frequencies. The resulting operation infidelity, $I$, is evaluated, to leading order in the qubit-qubit coupling deviations and in residual ion displacements, as,
@@
Figure 7(top) presents the infidelity of the initial LSF solution, based on the typical cell (orange) and the crosstalk corrected solution (dark-blue). It exhibits an apparent cross-over between two regimes at $\omega_{\text{o.t.p}}=2\nu$, at which point the transverse mode band structure is manifested. At low optical confinement the performance of uncorrected, typical-cell solutions (orange) is low, such that our leading-order estimation breaks-down. At strong optical confinement this solution generates an infidelity of $\sim 10^{-2}$, as expected from our analysis above. However the crosstalk-corrected solution (dark-blue) successfully generates solutions that have an infidelity that is better (lower) than $10^{-4}$ for the entire multi-qubit operation, throughout the strong optical confinement regime.
We account for the different contributions to the infidelity of the crosstalk-corrected solution, given by Eq. (7). We note that $I_{\text{short}}$ (dashed pink), has a negligible contribution, regardless of optical confinement, while $I_{\text{long}}$ (dashed purple) reaches a low values only at the strong confinement regime. This behavior is expected as our optimization actively corrects for on-site and nearest-neighbor terms, and ignores other terms. We note that $I_{\text{long}}$ reaches a minimum at $\omega_{\text{o.t.p}}=2\nu$, and is slightly increased at higher optical trapping frequencies. This behavior is accounted by an effective screening of barrier ions that occurs when barrier bands are close to resonance with the bulk bands (see the SM [48]). Nevertheless, our analysis implies that strong optical confinement successfully decouples next nearest-neighbor (and further) cells, removing the need to take the entire ion-crystal into account, and enabling scalability. Similarly, due to a well-formed band structure, the motional infidelity, $I_{\text{motion}}$ (dashed light-blue), becomes small at strong optical confinement, and in general dominates the overall infidelity.
@@
We remark that elements of our architecture and crosstalk compensation method can also be implemented by global driving of entire cells using similar techniques to those in Refs. [13, 15]. Moreover by employing the established methods in Refs. [70, 71, 72, 73, 74, 75, 76], we can impose additional robustness constraints on each gate drive in order to reduce the sensitivity of the gate and the crosstalk compensation, to fluctuations in trap parameters as well as other sources of noise.
@@ VII Mid-circuit measurements
Here however, optically confined ions are well-suited for mid-circuit measurements, as their motion is separated to independent bands (top part of the spectra in Figs. 4 and 5), thus photon scattering heating remains local and does not heat the other crystal modes. Furthermore, the optical trapping also substantially light-shifts the atomic transition lines of the ions, specifically the photon emission lines, such that secondary photon scattering from neighboring computational qubit ions is largely suppressed. Thus, no physical shuttling of the ions into dedicated measurement regions is required, nor the use of multiple ion-species.
Nevertheless, measured ions may heat substantially, thus each measurement step is followed by mid-circuit cooling. Only the modes of optically confined ions must be cooled this way, leaving the information encoded in computational qubit ions unaffected. Mid-circuit cooling can be done on all optically confined ions in parallel, and is expected to occupy a minor part of the overall circuit duration (see the SM for further detail [48]).
@@
Figure 8: Top: Schematic of mid-circuit measurement protocol, showing three types of involved ions: computational qubit ions (top, white circle), barrier ions ($B_{\text{B}}$) to be measured (middle, white circle), and barrier ions ($B_{\text{A}}$) that are currently optically confined (bottom, purple filled circle). Operations are ordered from left to right. Data is encoded on the computational qubit (e.g. $U_{s,n}^{\left(A\right)}$), and one of the qubit states of the ion is ‘shelved’ to a non-fluorescing state using local control. Next an intermediate optical segmentation configuration is used, confining both A and B barrier ions, and separating their motion from the computational qubits. The state of the barrier ions is measured and then reset (‘Prepare’). Lastly the optical segmentation configuration is switched. Bottom: Relevant atomic levels of ${}^{40}\text{Ca}^{+}$ ions, showing $4S_{1/2}$ the ground state manifold, the metastable $4D_{5/2}$ manifold and the short-lived $4P_{1/2}$ manifold. The $S\leftrightarrow P$ transition varies between the no-optical confinement (left) and optically confined (right) cases, allowing to separate the two cases spectrally (see main text). The main optical field used in the protocol are shown, namely at $729$ nm (red), coupling the $S\leftrightarrow D$ levels, at $400$ nm (light-blue), generating Raman transition between qubit states as well as additional local light-shifts, and at $397$ nm (purple) coupling the $S\leftrightarrow P$ levels.
We present mid-circuit measurements in conjunction with the ion-crystal’s segmentation reconfiguration, as the two processes naturally combine, though each of these can also be performed independently. The protocol consists of measuring the qubit states of some target ions, without incurring decoherence on computational qubits, or heating of bulk modes, after which we cool, re-initialize the measured ions and reconfigure the ion-crystal’s segmentation.
@@
Figure 8 (top) illustrates these general steps involved in performing mid-circuit measurements. For simplicity we consider the required operations on three types of ions in the ion-crystal, i.e. computational qubit ions (top, white circles), barrier ions to be measured (middle, white circle), and optically confined barrier ions (bottom, purple filled circles). The measurement method can be performed in parallel throughout the entire ion-crystal on all required ions according to these roles. In our architecture ions may be optically confined specifically for the purpose of mid-circuit measurements, without necessarily having an essential role in segmentation of the ion-crystal, as showcased in the application of a QEC surface code, above.
@@
An additional locally-applied field resonant with the light-shifted $S\leftrightarrow P$ transition at $397$ nm is used (purple arrow). Since the $P$ level is broad, the required stability of the optically confining fields in order to maintain resonance is attainable. The combination of local $400$ nm and $397$ nm fields, together with standard repump fields (not shown) accommodates for all the necessary steps - i.e., state detection, various cooling techniques, and state preparation. We note that the technical requirements for the $397$ nm field are less demanding than those of the optical confinement and gate drive fields at $400$ nm; the necessary power per ion is far lower, independent spectral control is not required, and the focal spot can generally be larger (by illuminating adjacent ions within the same beam).
@@
The mid-circuit measurement protocol shown in Fig. 8 is carried out by first encoding some information on ion to be measured (middle) via the entanglement operations detailed above ($U_{s,n}^{\left(A\right)}$), e.g. parity information of a plaquette shown in Fig. 3. Then, one of the qubit states of the ion to be measured is ‘shelved’ to the D manifold via a combination of $400$ nm light-shift and $729$ nm coupling, making this operation local on the ion.
Next, the segmentation configuration is changed to an intermediate setting where all barrier ions are illuminated. This separates the motion of measured ions from the computational qubits’ motion and light-shifts their $S\leftrightarrow P$ transition. At this step, some ions may fully populate the $D$ level, in which case they are not optically confined; thus existing confinement of other ions (bottom) cannot be relaxed.
@@
State detection is then performed using the light-shifted $397$ nm field, followed by qubit reset and preparation using a combination of the $397$ nm and $400$ nm fields. During the detection and preparation steps, the optically confined ions scatter $397$ nm photons. The scattered photons are substantially detuned, by approximately $0.8$ GHz, from the absorption lines of the computational qubit ions, having a linewidth of few tens of MHz, thus negligibly influencing their state. Nevertheless the state of neighboring computational qubits, located several micrometers apart, may be further protected by shelving both of their qubit states to the $D$ level (at the shelving step).
The measured ions are required to be ground state cooled in the ‘prepare’ step, which can be done on all optically confined ions in parallel. Incidentally, this cooling step also dissipates the heat that accumulates on barrier ions during the circuit due to off-resonant photon scattering and due to measurements.
During exposure to $397$ nm light, the measured ions inevitably spend some portion of time in the excited $P$ state, in which the optical potential is anti-trapping. Moreover, transitions to the $P$ state may introduce heating due to the fluctuating dipole force [84, 85]. The heat induced by these processes yields an excitation of $\sim 5$ motional quanta of their localized mode, keeping them in a manageable regime, from which they can be recooled back to their motional ground state, detailed in the SM [48].
Lastly, the segmentation configuration can be changed, switching the roles of computational and barrier ions. Although the barrier ions are cooled before returning to the bulk, care must be taken to reconfigure the optical potentials slowly enough so as not to excite any motion of the bulk modes. As detailed in the SM [48], this condition is satisfied as long as the reconfiguration time is large compared to an oscillation period of the ion trap; in practice, a reconfiguration time of $\sim 30\mu\text{s}$ easily meets this requirement. Overall the duration of the mid-circuit measurement, preparation, and reconfiguration steps combined is generally small compared to the multi-qubit gate time.